For a reaction A + 3B gives 2C + D initial mole of A is twice that of B. If at equilibrium moles of B and C are equal, then percent of B reacted is ?

Grade:12

1 Answers

Vikas TU
7 years ago
                   A + 3B --------> 2C + D t=0           2a     a                 0        0
t=t              2a-x   a-x               2x        x
Now given,
a-x = 2x = > a = 3x = > x = a/3.
[B] = 2a/3
[A] = 5a/3
% of B reacted is: => a/3a * 100 = 33.33 %